# Variational quantum state preparation for quantum-enhanced metrology in noisy systems

### v1

@@ III Results
In this section, we present results of our numerical optimization for the range of entangling Hamiltonians defined in Eqs. (2)-(4) and for extreme cases of $\alpha=0$ and $6$. A key result of our investigation is that we observe that the optimal states prepared by the VQC can be typically delineated into three regimes: (i) perturbative, characterized by weak decoherence such that the optimal preparation strategy generates macroscopic superposition states that approach the cat state that is known to achieve sensitivity at the HL in the absence of noise, (ii) squeezed-like, characterized by states with non-zero polarization (i.e., collective spin projection) and anisotropic projection noise, and (iii) uncorrelated, corresponding to a separable coherent spin state of the qubits with sensitivity near the SQL. This classification is based on analysis of the optimal QFI and abrupt transitions in the associated VQC parameters, but is further supported through the examination of a number of auxiliary quantities, including state overlap, distribution functions and spin squeezing, which provide additional information about the structure of the prepared quantum states.
@@
Figure 2: The normalized QFI $F_{Q}/N^{2}$ and optimal parameters $\mathbf{x}_{\mathrm{opt}}$ as functions of isotropic decoherence strength, $\gamma$ for (a) OAT and TAT, and (b) Ising, XX and f-TAT models. In all panels we use $N=8$ and the interaction range is set to (a) $\alpha=0$ and (b) $\alpha=6$. For all cases we exclude the initial rotation angle $\theta_{1}$, as it is found to identically be either $\pi/2$ (OAT, Ising and XX) or 00 (TAT and f-TAT) regardless for non-trivial values of $\gamma<\gamma^{(2)}$ (see main text). The values of $\gamma^{(1)}$ and $\gamma^{(2)}$ for each model are indicated by vertical lines (blue dashed for OAT and Ising, red dot-dashed for TAT and f-TAT and solid green for XX) to guide the reader between the main regimes. Note that $\gamma^{(1)}$ is absent for the Ising data, while the values of $\gamma^{(1)}$ for XX and f-TAT are indistinguishable.
III.1 One- and two-axis twisting
@@
We begin by presenting results for OAT and TAT with infinite-range interactions ($\alpha=0$), as the dynamics generated by these Hamiltonians is well understood in the absence of noise Kitagawa and Ueda (1993); Kajtoch and Witkowska (2015) and they can provide a template for our later discussion of finite-range interactions. The optimal QFI and parameters obtained for OAT and TAT are plotted in panel (a) of Fig. 2 as a function of the isotropic dephasing strength $\gamma$ and with fixed $N=8$. For both cases, the QFI begins near the maximal value $F_{Q}/N^{2}\approx 1$ (HL) for $\gamma\ll 1$ (in fact, it is essentially saturated for OAT) and then monotonically decreases until settling at the SQL $F_{Q}/N^{2}=1/N$ for $\gamma\gg 1$.
@@
We observe that the QFI appears to be continuous as function of $\gamma$ for both OAT and TAT, although there is a clear kink in the QFI at $\gamma\approx 0.25$ for OAT. On the other hand, the associated VQC parameters $\mathbf{x}_{\mathrm{opt}}$ show more abrupt behaviour, including discontinuities, that we use to qualitatively classify different regimes of quantum states. In particular, we use the strength of the entangling gate $\theta_{\mathrm{I}}$ to define two characteristic values of $\gamma$. For vanishingly small $\gamma$ we observe that $\theta_{I}$ takes the largest values for both OAT and TAT, but it slowly decreases with increasing $\gamma$. We identify $\gamma^{(1)}$ as the dephasing strength at which $\theta_{I}$ exhibits a discontinuity and abruptly drops in value, leading to $\gamma^{(1)}\approx 0.25$ and $0.02$ for OAT and TAT, respectively. Further decrease in $\theta_{I}$ is then observed for $\gamma>\gamma^{(1)}$ until it vanishes, i.e., the optimal VQC does not feature any application of the entangling gate. We define the point at which $\theta_{I}=0$ as $\gamma^{(2)}$, and find $\gamma^{(2)}\approx 9.44$ and $25.85$ for OAT and TAT, respectively. Additionally, we observe that similar discontinuities or abrupt behaviour are present in other VQC parameters, e.g., rotation angle $\theta_{2}$ and the axis difference $|\varphi_{1}-\varphi_{2}|/\pi$. Note that we do not show data for the initial rotation angle $\theta_{1}$, as we find that the optimal choice is $\theta_{1}=\pi/2$ for OAT and $\theta_{1}=0$ for TAT for $\gamma<\gamma^{(2)}$ (i.e., “non-trivial” values of $\gamma$).
@@
Insight into the three distinct parameter regimes carved out by the definition of $\gamma^{(1)}$ and $\gamma^{(2)}$ is found by drilling further into the OAT results, which can be contrasted against the well-understood unitary limit of $\gamma=0$. Firstly, when $\gamma<\gamma^{(1)}$ we observe that the optimal parameters $\mathbf{x}_{\mathrm{opt}}$ [see Fig. 2 panel (a)] are virtually identical to those that would produce a GHZ state [see Eq. (7)] in the absence of decoherence Foss-Feig et al. (2013). In particular, our optimal VQC sequence corresponds to a first rotation of $\theta_{1}=\pi/2$ with arbitrary value of $\varphi_{1}$, which orients the initially prepared product state on the collective Bloch sphere equator. The rotation is followed by an entangling gate with strength $\theta_{\mathrm{I}}\approx\pi/2$, which is consistent with the value required to generate a GHZ state using twisting dynamics Foss-Feig et al. (2013). At this point, the generated GHZ-like state is characterized by a collective superposition of qubits oppositely polarized along an axis lying in the equatorial plane (specifically, along an axis perpendicular to that defined by the first rotation about $\varphi_{1}$). Thus, the final rotation by an angle $\theta_{2}=\pi/2$ about the axis $\varphi_{2}\approx\varphi_{1}$ reorients the GHZ-like state along the $z$ axis [similar to Eq. (7)]. In the absence of noise, this state features the largest possible fluctuations in the observable $\hat{S}_{z}$ and thus maximizes the QFI Tóth and Apellaniz (2014). As a result, we dub $\gamma<\gamma^{(1)}$ as the perturbative regime, as our results suggest that decoherence is sufficiently weak so that the optimal state preparation strategy is simply to closely (i.e., perturbatively) follow the approach established in the absence of decoherence.
@@
On the other hand, in the regime where $\gamma^{(1)}<\gamma<\gamma^{(2)}$, $\theta_{\mathrm{I}}$ takes on relatively smaller values that can be associated with the paradigmatic generation of spin-squeezed states Kitagawa and Ueda (1993). Particular to one-axis twisting, the unitary generation of spin-squeezed stated occurs via “shearing” of the quantum fluctuations of an initial coherent spin state polarized along the equator of the collective Bloch sphere. The resulting squeezing of the projection noise is thus in a spin quadrature that is rotationally offset from the equatorial plane of the Bloch sphere, by an amount that depends on the strength/duration of the twisting interaction ($\theta_{I}$), and thus requires a final rotation about an axis aligned along the collective polarization of the state. We see signatures of these features in our dynamics with decoherence, including that the final rotation axis is perpendicular to the initial rotation (and thus parallel to the polarization of the state post first rotation), $|\varphi_{2}-\varphi_{1}|\approx\pi/2$, whilst the angle of the rotation $\theta_{2}$ varies with the duration of the entangling interaction characterized by $\theta_{\mathrm{I}}$ Kitagawa and Ueda (1993). Thus, since our results suggest that the states within this regime may be closely associated with spin-squeezing we designate $\gamma^{(1)}<\gamma<\gamma^{(2)}$ as the squeezed-like regime.
@@
Finally, we denote $\gamma>\gamma^{(2)}$ as the uncorrelated regime. In this case, as the entangling gate is never applied ($\theta_{I}=0$), the VQC reduces to a pair of global rotations that together simply reorient the initial pure, uncorrelated coherent spin state onto some axis in the $x$-$y$ plane to maximize the QFI at the SQL, $F_{Q}/N^{2}=1/N$. This regime arises for overwhelmingly large decoherence, such that any application of the entangling gate would in fact lead to a QFI below the SQL due to an increase in total projection noise, i.e., the interactions cannot generate metrologically useful correlations faster than decoherence degrades the quantum state [see the bottom subplot of Fig. 3].
@@
Figure 3: Diagnostic quantities calculated for optimally prepared states as functions of dephasing strength $\gamma$ for (a) OAT and (b) TAT. All data is for $N=8$. (Top subplots) Normalized classical Fisher information ($F_{c}/N^{2}$, squares), inverse Wineland squeezing parameter $(N\xi^{2}_{s})^{-1}$, tri-prongs) and the normalized QFI ($F_{Q}/N^{2}$, circles for OAT and triangles for TAT). (Lower subplots) The overlap of the optimal state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$, diamonds) and the normalized collectivity ($\langle\hat{S}^{2}\rangle/S(S+1)$, crosses). Vertical lines in all panels indicate $\gamma^{(1)}$ and $\gamma^{(2)}$. Below each set of panels we show Husimi phase space distributions for representative quantum states in each regime (see Appendix A for details of distribution). Bright regions indicate larger probability density. For the GHZ-like states we show two perspectives of the same quantum state, rotated by $\pi/2$ about indicated axes.
@@
The behaviour of the optimal parameters $\mathbf{x}_{\mathrm{opt}}$ for TAT (also shown in Fig. 2) follow similar trends and is suggestive of a similar classification into three state preparation regimes. However, in this case the increased complexity of the unitary dynamics for TAT Kajtoch and Witkowska (2015) makes quantitative analysis of the parameters more difficult. Thus, in the following we build on the intuition developed from OAT and perform a more systematic investigation of the structure of the prepared quantum states using a variety of diagnostic quantities.
III.1.1 Characterization of GHZ-like states
@@
We first focus on the perturbative regime defined by $\gamma<\gamma^{(1)}$. Motivated by the observation that the QFI in this regime is near the HL ($F_{Q}/N^{2}=1$) and informed by the close proximity of the parameters $\mathbf{x}_{\mathrm{opt}}$ for OAT to those that are known to generate GHZ states [see Eq. (7)] for $\gamma=0$, we calculate the overlap between the optimally prepared state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ and a GHZ state. Specifically, we compute the maximum fidelity,
@@
The results for this quantity as a function of $\gamma$ are shown in the middle panels of Fig. 3(a) and (b). We observe appreciable fidelities $1/2<\mathcal{F}_{\mathrm{GHZ}}\leq 1$ in the regime $\gamma<\gamma^{(1)}$ for both OAT [panel (a)] and TAT [panel (b)]. The values for the fidelity are indicative of strong overlap with the GHZ state and the presence of $N$-qubit entanglement Leibfried et al. (2005). For both OAT and TAT the fidelity remains appreciable when $\gamma>\gamma^{(1)}$, although it steadily decreases Huelga et al. (1997). In particular, the fidelity appears to be more robust to increasing decoherence for TAT. However, we qualify these observations by noting that the system size used in our calculations is quite small.
@@ III.1.2 Characterization of squeezed-like states
In the squeezed-like regime found for larger decoherence, $\gamma^{(1)}<\gamma<\gamma^{(2)}$, we expect the prepared states to have a non-zero collective polarization and anisotropic quantum fluctuations in the spin quadratures defined perpendicular to the orientation of the state. Mathematically, this structure can be captured by the Wineland squeezing parameter Wineland et al. (1992),
@@
where $\langle\hat{\mathbf{S}}\rangle=(\langle\hat{S}_{x}\rangle,\langle\>\hat{S}_{y}\rangle,\>\langle\hat{S}_{z}\rangle)$ is the collective Bloch vector of the state and $\mathbf{n}_{\perp}$ denotes a unit vector lying in a plane orthogonal to $\langle\hat{\mathbf{S}}\rangle$. The squeezing parameter is obtained by minimizing the fluctuations $\langle(\Delta\hat{S}_{\mathbf{n}_{\perp}})^{2}\rangle$ with respect to the orientation of $\mathbf{n}_{\perp}$. A value $\xi^{2}_{s}<1$ reflects that the quantum fluctuations along one spin quadrature are squeezed below the projection noise limit such that the state can enable sensing of spin rotations below the SQL using a simple Ramsey sequence Wineland et al. (1992). More generally, squeezing sets a lower bound on the QFI (and thus metrological utility of the quantum state) through the inequality $F_{Q}/N^{2}\geq(N\xi^{2}_{s})^{-1}$.
We plot the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$ as a function of $\gamma$ in the upper panels of Fig. 3(a) and (b). We observe that $(N\xi^{2}_{s})^{-1}=0$ in the perturbative regime of $\gamma<\gamma^{(1)}$ as the polarization of the cat state vanishes, i.e., $\langle\hat{\mathbf{S}}\rangle=0$. On the other hand, for $\gamma>\gamma^{(1)}$ we observe $(N\xi^{2}_{s})^{-1}>0$. Initially, as $\gamma$ is increased past this transition, the inverse squeezing parameter does not breach the SQL, i.e., $(N\xi^{2}_{s})^{-1}<1/N$, indicating that while the QFI is large ($F_{Q}/N^{2}>1/N$) and corresponds to a quantum-enhanced sensitivity, the state does not exhibit simple squeezing that can be exploited using Ramsey interferometry with mean values of global spin measurements. We attribute this to oversqueezing, which has been studied in the context of OAT dynamics Strobel et al. (2014); Bohnet et al. (2016). To demonstrate this, we show example Husimi distributions for representative quantum states in this regime below panels Fig. 3(a) and (b). These clearly illustrate that the quantum state features anisotropic quantum noise. However, the distribution wraps appreciably around the Bloch sphere, leading to a shrinking of the collective Bloch vector $\langle\hat{\mathbf{S}}\rangle$ and the failure of the squeezing parameter to capture the metrological sensitivity of the state. As $\gamma$ further increases, the squeezing parameter breaks through the SQL and eventually matches the QFI, $(N\xi^{2}_{s})^{-1}=F_{Q}/N^{2}$. In this latter regime, the metrological utility of the state is thus entirely captured by spin squeezing.

### v2

@@ III Optimal state preparation with isotropic dephasing
In this section we present the results of our numerical optimization for the range of entangling Hamiltonians defined in Eqs. (2)–(4). Moreover, we focus on $\alpha=0$, $3$ and $6$ and isotropic dephasing noise such that $\gamma_{x}=\gamma_{y}=\gamma_{z}=\gamma$. This assumption voids bias in our results due to, for example, robustness of the entangling dynamics for specific representations of the Hamiltonians in Eqs. (2)–(4Britton et al. (2012); Jurcevic et al. (2014). A key result of our investigation is that we observe that the states prepared by the optimal VQC can be typically delineated into three regimes: (i) a cat-like regime, characterized by quantum states typified by Eq. (7) but more broadly described as a superposition of macroscopically distinct projections, (ii) a squeezed-like regime, characterized by states with non-zero polarization (i.e., collective spin projection) and anisotropic projection noise, and (iii) an uncorrelated regime, corresponding to a separable coherent spin state of the qubits with sensitivity near the SQL. This classification is based on an analysis of the optimal QFI and abrupt transitions in the associated VQC parameters but is further supported through the examination of auxiliary quantities, including state overlap, distribution functions, and spin squeezing, which provide additional information about the structure of the prepared quantum states.
@@
Figure 2: The normalized QFI $F_{Q}/N^{2}$ and optimal VQC parameters $\mathbf{x}_{\mathrm{opt}}$ as functions of isotropic decoherence strength, $\gamma$ (in units of $\chi$) for (a) OAT and TAT and for (b) and (c) Ising, XX, and f-TAT models. In all panels we use $N=8$, and the interaction range is set to (a) $\alpha=0$, (b) $\alpha=3$, and (c) $\alpha=6$. For all cases we exclude the initial rotation angle $\theta_{1}$, since it is found to identically be either $\pi/2$ (OAT, Ising, and XX) or 00 (TAT and f-TAT) regardless for nontrivial values of $\gamma<\gamma^{(2)}$ (see main text). The values of $\gamma^{(1)}$ and $\gamma^{(2)}$ for each model are indicated by vertical lines (blue dashed for OAT and Ising, red dot-dashed for TAT and f-TAT, and solid green for XX) to guide the reader between the main regimes. Note that $\gamma^{(1)}$ is absent for the Ising data, while the values of $\gamma^{(1)}$ for XX and f-TAT in panel (c) are indistinguishable.
III.1 Infinite-range interactions: One- and two-axis twisting
@@
We begin by presenting results for OAT and TAT with infinite-range interactions ($\alpha=0$), since the dynamics generated by these Hamiltonians is well understood in the absence of noise Kitagawa and Ueda (1993); Kajtoch and Witkowska (2015) and they can provide a template for our later discussion of finite-range interactions. The optimal QFI and parameters obtained for OAT and TAT are plotted in panel (a) of Fig. 2 as a function of the isotropic dephasing strength $\gamma$ (in units of $\chi$) and with fixed $N=8$. For both cases, the QFI begins near the maximal value $F_{Q}/N^{2}\approx 1$ (HL) for $\gamma/\chi\ll 1$ (in fact, it is essentially saturated for OAT) and then monotonically decreases until settling at the SQL ($F_{Q}/N^{2}=1/N$) for $\gamma/\chi\gg 1$.
@@
We observe that the QFI appears to be continuous as a function of $\gamma$ for both OAT and TAT, although there is a clear kink in the QFI at $\gamma/\chi\approx 0.25$ for OAT. On the other hand, the associated VQC parameters $\mathbf{x}_{\mathrm{opt}}$ show more abrupt behavior, including discontinuities, that we use to qualitatively classify different regimes of quantum states. In particular, we use the strength of the entangling gate $\theta_{\mathrm{I}}$ to define two characteristic values of $\gamma$. For vanishingly small $\gamma$ we observe that $\theta_{I}$ takes the largest values ($\theta_{\mathrm{I}}/\pi\approx 0.5$) for both OAT and TAT, but it decreases with increasing $\gamma$. We identify $\gamma^{(1)}$ as the dephasing strength at which $\theta_{I}$ exhibits a discontinuity and abruptly drops in value, leading to $\gamma^{(1)}/\chi\approx 0.25$ and $0.02$ for OAT and TAT, respectively. Further decrease in $\theta_{I}$ is then observed for $\gamma>\gamma^{(1)}$ until it vanishes corresponding to turning off the entangling gate. We define the point at which $\theta_{I}=0$ as $\gamma^{(2)}$ and find $\gamma^{(2)}/\chi\approx 9.44$ and $25.85$ for OAT and TAT, respectively. Additionally, we observe that similar discontinuities or abrupt behavior are present in other VQC parameters, for example, rotation angle $\theta_{2}$ and the axis difference $|\varphi_{1}-\varphi_{2}|/\pi$. Note that we do not show data for the initial rotation angle $\theta_{1}$, since we find that the optimal choice is $\theta_{1}=\pi/2$ for OAT and $\theta_{1}=0$ for TAT for $\gamma<\gamma^{(2)}$ (i.e., “nontrivial” values of $\gamma$).
@@
Insight into the three distinct parameter regimes carved out by the definition of $\gamma^{(1)}$ and $\gamma^{(2)}$ is found by drilling further into the OAT results, which can be contrasted against the well-understood unitary limit of $\gamma/\chi=0$. First, when $\gamma<\gamma^{(1)}$, we observe that the optimal parameters $\mathbf{x}_{\mathrm{opt}}$ [see Fig. 2 panel (a)] are virtually identical to those that would produce a GHZ state [see Eq. (7)] in the absence of decoherence Foss-Feig et al. (2013). In particular, our optimal VQC sequence corresponds to a first rotation of $\theta_{1}=\pi/2$ with an arbitrary value of $\varphi_{1}$, which orients the initially prepared product state on the collective Bloch sphere equator. The rotation is followed by an entangling gate with strength $\theta_{\mathrm{I}}\approx\pi/2$, which is consistent with the value required to generate a GHZ state using twisting dynamics Foss-Feig et al. (2013). At this point, the generated GHZ-like state is characterized by a collective superposition of qubits oppositely polarized along an axis lying in the equatorial plane (specifically, along an axis perpendicular to that defined by the first rotation about $\varphi_{1}$). Thus, the final rotation by an angle $\theta_{2}=\pi/2$ about the axis $\varphi_{2}\approx\varphi_{1}$ reorients the GHZ-like state along the $z$-axis [similar to Eq. (7)]. In the absence of noise, this state features the largest possible fluctuations in the observable $\hat{S}_{z}$ and thus maximizes the QFI Tóth and Apellaniz (2014). As a result, we dub $\gamma<\gamma^{(1)}$ as the cat-like regime, since our results suggest that decoherence is sufficiently weak that the optimal state preparation strategy is simply to closely (i.e., cat-likely) follow the approach established in the absence of decoherence.
@@
On the other hand, in the regime where $\gamma^{(1)}\leq\gamma<\gamma^{(2)}$, $\theta_{\mathrm{I}}$ takes on relatively smaller values that can be associated with the paradigmatic generation of spin-squeezed states Kitagawa and Ueda (1993). Particular to one-axis twisting, the unitary generation of spin-squeezed states occurs via “shearing” of the quantum fluctuations of an initial coherent spin state polarized along the equator of the collective Bloch sphere. The resulting squeezing of the projection noise is thus in a spin quadrature that is rotationally offset from the equatorial plane of the Bloch sphere, by an amount that depends on the strength/duration of the twisting interaction ($\theta_{I}$), and thus requires a final rotation about an axis aligned along the collective polarization of the state. We see signatures of these features in our dynamics with decoherence, including that the final rotation axis is perpendicular to the initial rotation (and thus parallel to the polarization of the state post first rotation), $|\varphi_{2}-\varphi_{1}|\approx\pi/2$, while the angle of the rotation $\theta_{2}$ varies with the duration of the entangling interaction characterized by $\theta_{\mathrm{I}}$ Kitagawa and Ueda (1993). Thus, since our results suggest that the states within this regime may be closely associated with spin squeezing, we designate $\gamma^{(1)}\leq\gamma<\gamma^{(2)}$ as the squeezed-like regime.
@@
Finally, we denote $\gamma\geq\gamma^{(2)}$ as the uncorrelated regime. In this case, since the entangling gate is never applied ($\theta_{I}=0$), the VQC reduces to a pair of global rotations that together simply reorient the initial pure, uncorrelated coherent spin state onto some axis in the $x$-$y$ plane to maximize the QFI at the SQL, $F_{Q}/N^{2}=1/N$. This regime arises for overwhelmingly large decoherence, such that any application of the entangling gate would in fact lead to a QFI below the SQL due to an increase in total projection noise. In other words, the interactions cannot generate metrologically useful correlations faster than decoherence degrades the quantum state [see the bottom subplot of Fig. 3].
@@
Figure 3: Diagnostic quantities calculated for optimally prepared states as functions of isotropic dephasing strength $\gamma$ (in units of $\chi$) for (a) OAT and (b) TAT (i.e., Ising and f-TAT Hamiltonians, respectively, with $\alpha=0$). All data is for $N=8$. Top subplots: Normalized classical Fisher information ($F_{c}/N^{2}$, squares), inverse Wineland squeezing parameter $(N\xi^{2}_{s})^{-1}$, tri-prongs) and the normalized QFI ($F_{Q}/N^{2}$, circles for OAT and triangles for TAT). Lower subplots: Overlap of the optimal state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ with a GHZ state ($\mathcal{F}_{\mathrm{GHZ}}$, diamonds) and the normalized collectivity ($\langle\hat{S}^{2}\rangle/S^{2}_{\mathrm{max}}$, crosses). Vertical lines in all panels indicate $\gamma^{(1)}$ and $\gamma^{(2)}$. Below each set of panels we show Husimi phase space distributions for representative quantum states in each regime (see Appendix A for details of distribution). Bright regions indicate larger probability density. For $\gamma/\chi=0.01$ we show two perspectives of the same quantum state, rotated for indicated axes and angles.
@@
The behavior of the optimal parameters $\mathbf{x}_{\mathrm{opt}}$ for TAT (also shown in Fig. 2) follows similar trends and is suggestive of a similar classification into three state preparation regimes. In this case, however, the increased complexity of the unitary dynamics for TAT Kajtoch and Witkowska (2015) makes quantitative analysis of the parameters more difficult. Thus, in the following we build on the intuition developed from OAT and perform a more systematic investigation of the structure of the prepared quantum states using a variety of diagnostic quantities.
III.1.1 Characterization of cat-like states
@@
We first focus on the cat-like regime defined by $\gamma<\gamma^{(1)}$. Motivated by the observation that the QFI in this regime is near the HL ($F_{Q}/N^{2}=1$) and informed by the close proximity of the parameters $\mathbf{x}_{\mathrm{opt}}$ for OAT to those that are known to generate GHZ states [see Eq. (7)] for $\gamma/\chi=0$, we calculate the overlap between the optimally prepared state $\hat{\rho}(\mathbf{x}_{\mathrm{opt}})$ and a GHZ state. Specifically, we compute the maximum fidelity,
@@
The results for this quantity as a function of $\gamma/\chi$ are shown in the middle panels of Fig. 3(a) and (b). We observe appreciable fidelities $1/2<\mathcal{F}_{\mathrm{GHZ}}\leq 1$ in the regime $\gamma<\gamma^{(1)}$ for both OAT [panel (a)] and TAT [panel (b)]. The values for the fidelity are indicative of strong overlap with the GHZ state and the presence of $N$-qubit entanglement Leibfried et al. (2005). Relative to the OAT results, the TAT case is observed to typically feature a reduced fidelity to the GHZ state, indicating that an ideal GHZ is never formed even for $\gamma/\chi\to 0$. Interesting, however, $\mathcal{F}_{\mathrm{GHZ}}$ remains robustly above or quite near $1/2$ even for some values of $\gamma>\gamma^{(1)}$, although it steadily decreases. We attribute this to the inherent richness of states that can generated by TAT interactions Kajtoch and Witkowska (2015) and the relatively faster generation of entanglement that is expected Kitagawa and Ueda (1993). In contrast, the fidelity for OAT rapidly falls below $1/2$ prior to $\gamma=\gamma^{(1)}$. However, we qualify these observations by noting that the system size used in our calculations is quite small.
@@ III.1.2 Characterization of squeezed-like states
In the squeezed-like regime found for larger decoherence, $\gamma^{(1)}\leq\gamma<\gamma^{(2)}$, we expect the prepared states to have a non-zero collective polarization and anisotropic quantum fluctuations in the spin quadratures defined perpendicular to the orientation of the state. Mathematically, this structure can be captured by the Wineland squeezing parameter Wineland et al. (1992),
@@
where $\langle\hat{\mathbf{S}}\rangle=(\langle\hat{S}_{x}\rangle,\langle\>\hat{S}_{y}\rangle,\>\langle\hat{S}_{z}\rangle)$ is the collective Bloch vector of the state and $\mathbf{n}_{\perp}$ denotes a unit vector lying in a plane orthogonal to $\langle\hat{\mathbf{S}}\rangle$. The squeezing parameter is obtained by minimizing the fluctuations $\langle(\Delta\hat{S}_{\mathbf{n}_{\perp}})^{2}\rangle$ with respect to the orientation of $\mathbf{n}_{\perp}$. A value $\xi^{2}_{s}<1$ reflects that the quantum fluctuations along one spin quadrature are squeezed below the projection noise limit such that the state can enable sensing of spin rotations below the SQL using a simple Ramsey sequence Wineland et al. (1992). More generally, squeezing sets a lower bound on the QFI (and thus the metrological utility of the quantum state) through the inequality $F_{Q}/N^{2}\geq(N\xi^{2}_{s})^{-1}$.
We plot the inverse squeezing parameter $(N\xi^{2}_{s})^{-1}$ as a function of $\gamma$ in the upper panels of Fig. 3(a) and (b). We observe that $(N\xi^{2}_{s})^{-1}=0$ in the cat-like regime of $\gamma<\gamma^{(1)}$ as the polarization of the cat state vanishes; that is, $\langle\hat{\mathbf{S}}\rangle=0$. On the other hand, for $\gamma\geq\gamma^{(1)}$ we observe $(N\xi^{2}_{s})^{-1}>0$. Initially, as $\gamma$ is increased past this transition, the inverse squeezing parameter does not breach the SQL, that is, $(N\xi^{2}_{s})^{-1}<1/N$, indicating that while the QFI is large ($F_{Q}/N^{2}>1/N$) and corresponds to a quantum-enhanced sensitivity, the state does not exhibit simple squeezing that can be exploited using Ramsey interferometry with mean values of global spin measurements. We attribute this to oversqueezing, which has been studied in the context of OAT dynamics Strobel et al. (2014); Bohnet et al. (2016). To demonstrate this, we show example Husimi distributions for representative quantum states in this regime below panels in Fig. 3(a) and (b). These clearly illustrate that the quantum state features anisotropic quantum noise. However, the distribution wraps appreciably around the Bloch sphere, leading to a shrinking of the collective Bloch vector $\langle\hat{\mathbf{S}}\rangle$ and the failure of the squeezing parameter to capture the metrological sensitivity of the state. As $\gamma$ further increases, the squeezing parameter breaks through the SQL and eventually matches the QFI, $(N\xi^{2}_{s})^{-1}=F_{Q}/N^{2}$. In this latter regime, the metrological utility of the state is thus entirely captured by spin squeezing.